### v1

# Qubit Mapping and Routing tailored to Advanced Quantum ISAs: Not as Costly as You Think

Abstract.
Qubit mapping/routing is a critical stage in compilation for both near-term and fault-tolerant quantum computers, yet existing scalable methods typically impose several times the routing overhead in terms of circuit depth or duration. This inefficiency stems from a fundamental disconnect: compilers rely on an abstract routing model (e.g., three-$\mathrm{CX}$-unrolled SWAP insertion) that completely ignores the idiosyncrasies of native gates supported by physical devices.
Recent hardware breakthroughs have enabled high-precision implementations of diverse instruction set architectures (ISAs) beyond standard $\mathrm{CX}$-based gates. Advanced ISAs involving gates such as $\mathrm{\sqrt{iSWAP}}$ and $\mathrm{ZZ}(\theta)$ gates offer superior circuit synthesis capabilities and can be realized with higher fidelities. However, systematic compiler optimization strategies tailored to these advanced ISAs are lacking.
To address this, we propose Canopus, a unified qubit mapping/routing framework applicable to diverse quantum ISAs. Built upon the canonical representation of two-qubit gates, Canopus centers on qubit routing to perform deep co-optimization in an ISA-aware approach. Canopus leverages the two-qubit canonical representation and the monodromy polytope to model the synthesis cost for more intelligent $\mathrm{SWAP}$ insertion during the routing stage. We also formalize the commutation relations between two-qubit gates through the canonical form, providing a generalized approach to commutativity-based optimizations. Experiments show that Canopus consistently reduces routing overhead by 15%-35% compared to state-of-the-art methods across different ISAs and topologies. Our work also presents a coherent method for co-exploration of program patterns, quantum ISAs, and hardware topologies. We have for the first time demonstrated that advanced quantum ISAs can be efficiently utilized within a unified routing framework, paving the way for more effective co-design of quantum software and hardware.
1. Introduction
Figure 1. Compilation workflows by means of conventional approaches (top) and Canopus (bottom) targeting diverse quantum ISAs. Canopus integrates the synthesis cost model (monodromy polytopes within the Weyl chamber) to consider backend ISA properties during the routing stage. Canopus routing operates in the 2Q canonical representation while the specific synthesis is completed by the backend synthesizer.
Quantum computing is a revolutionary computational paradigm leveraging quantum mechanical principles such as superposition and entanglement of qubit states (nielsen2010quantum). It has grown rapidly in recent decades due to the potential speedup in tasks such as integer factorization (shor1994algorithms), solving linear equations (harrow2009quantum), and simulation of quantum systems (lloyd1996universal).
Holistic benchmarks of quantum computers such as quantum volume (cross2019validating) are predicated on concurrent advancements in both hardware and software. Recently, numerous systematic techniques regarding compiler optimization and architectural support have been presented to push the limit of hardware performance. Quantum compilers play a pivotal role in this process, translating high-level programs into executable single-qubit (1Q) and two-qubit (2Q) gates on realistic quantum hardware. This typically involves several stages: (1) compiling programs into basic quantum gates, (2) performing hardware-agnostic (logical-level) circuit optimization, (3) resolving backend topology constraints via qubit placement and routing, and (4) converting circuits to native gates for final optimization and scheduling. The primary goal of compiler optimization is to lower the 2Q gate count and circuit depth while resolving backend constraints, with a particular emphasis on 2Q gates due to their significantly higher error rates compared to 1Q gates.
For mainstream quantum platforms such as superconducting qubits (krantz2019quantum), 2Q gates can only operate between nearest-neighbor physical qubit pairs (e.g., Google’s devices with 2D square topology (arute2019quantum), IBM’s devices with 2D heavy-hex topology (chamberland2020topological)). Consequently, qubit placement and routing is crucial for resolving this connectivity constraint by dynamically remapping logical qubits to physical ones by inserting $\mathrm{SWAP}$ gates acting on adjacent physical qubit pairs. This introduces a routing overhead that typically increases the gate count and circuit depth by a factor of 2x-4x relative to the pre-mapped circuits when using state-of-the-art (SOTA) scalable routing methods (li2019tackling; zulehner2018efficient; zhang2021time; liu2023tackling). Therefore, mitigating this routing overhead remains a central and long-standing challenge in compiler optimization.
Most studies on qubit routing rely on a simplified routing model, where circuit cost is quantified by the $\mathrm{CX}$-based gate count and circuit depth while each $\mathrm{SWAP}$ gate is unrolled into three $\mathrm{CX}$ gates according to the textbook pattern $\mathrm{SWAP}_{q_{0},q_{1}}=\mathrm{CX}_{q_{0},q_{1}}\mathrm{CX}_{q_{1},q_{0}}\mathrm{CX}_{q_{0},q_{1}}$. However, this $\mathrm{CX}$-centric view is misaligned with the physical reality of modern quantum devices. Although quantum algorithms are typically expressed in terms of $\mathrm{CX}$ gates, the underlying hardware may not execute native $\mathrm{CX}$-equivalent gates, nor does this gate/circuit cost quantification method accurately reflect the true operational cost. Indeed, beyond the native support for $\mathrm{CX}$-equivalent gates (e.g., $\mathrm{CZ}$ (krantz2019quantum), Cross-Resonance (rigetti2010fully), Mølmer-Sørensen (bruzewicz2019trapped)), modern quantum hardwares increasingly feature diverse native 2Q basis gates in recent years. These alternative basis gates, or the abstracted instruction set architectures (ISAs) can be more powerful than $\mathrm{CX}$-equivalent gates in terms of synthesis capabilities and fidelity of realization, such as $\mathrm{\sqrt{iSWAP}}$ (huang2023quantum), the $\mathrm{iSWAP}$-family and $\mathrm{CX}$-family gates (mckinney2024mirage; qiskitXXDecomposer), and heterogeneous/combinatorial basis gates (peterson2022optimal; mckinney2024mirage). With such ISAs, $\mathrm{SWAP}$ can be implemented with a lower cost than three $\mathrm{CX}$ gates or even be natively realized with high fidelity (wei2024native; chen2025efficient; nguyen2024programmable). Therefore, the simplified routing model completely ignores the backend ISA properties, severely limiting the potential of compiler optimization. Furthermore, the absence of systematic compiler optimization methods across these diverse (even complex, heterogeneous) ISAs has prevented the community from fully exploiting their power and exploring the rich software-hardware co-design space.
In our work, we propose a unified qubit mapping/routing framework Canopus (Canonical-Optimized Placement Utility Suite) tailored to diverse quantum ISAs. Unlike conventional $\mathrm{CX}$-based routing approaches, Canopus is fundamentally ISA-aware. As illustrated in Figure 1, it considers the properties of the target ISA by formulating an appropriate cost model to perform deep co-optimization of routing and synthesis. By means of the canonical 2Q representation (zhang2003geometric), Canopus fully exploits the synthesis capabilities of the given ISA. This approach demonstrates that advanced ISAs can achieve significantly lower routing overheads than conventional models suggest.
The main ideas of Canopus are as follows: ① Considering the subsequent synthesis overhead with native basis gates after the qubit routing stage could unlock significant optimization opportunities. For instance, a $\mathrm{SWAP}$ inserted after a 2Q block with the same qubit pair acted on could lead to a better synthesis result when rebased to the backend ISA. ② Expanding the quantum ISA is crucial to boost the performance of real-world quantum applications. For example, the fractional $\mathrm{ZZ}(\theta)$ gate set widely adopted by hardware vendors (e.g., IBM (ibmFractionalGates), Quantinuum (quantinuumArbitraryAngleGates), IonQ (ionqPartialGates)) enables more efficient execution of chemistry simulation kernels within which many 2-local Pauli rotations are involved. The combination of $\mathrm{CX}$ and $\mathrm{iSWAP}$ gates have been demonstrated to benefit stabilizer circuits to protect error-corrected qubit information (zhou2024halma). ③ The monodromy polytope theory (peterson2020fixed) based on the canonical representation of 2Q gates (zhang2003geometric) provides a formal, universal, and quantitative description of the 2Q synthesis cost for generic quantum ISAs. This formalism enables unified compiler optimization. With these insights in mind, Canopus performs intelligent $\mathrm{SWAP}$ insertion during qubit routing to holistically minimize post-mapping circuit cost (in terms of both gate count and depth) given any quantum ISA, thus performing deep routing-synthesis co-optimization and resulting in significantly lower routing overheads. Importantly, although Canopus is ISA-aware, it always operates on the canonical-form circuits, and the gate/circuit cost quantification via monodromy polytopes is independent of the specific ISA rebase implementation. Thus Canopus offers an LLVM-style optimization.
Experimental results demonstrate that Canopus consistently provides 15%-35% reduction (in terms of both gate count and depth) of routing overhead compared to other SOTA methods across representative quantum ISAs, including the conventional CX ISA. This cross-ISA comparison also reveals some consistent or program-specific or topology-specific guidelines for hardware-software co-design. Source code and data are available via the Github link. Our work makes the following key contributions:
❶We utilize the canonical 2Q gate representation and the monodrome polytope to quantify costs of 2Q gates and the overall circuit. This formal approach accurately guides synthesis-routing co-optimization and cross-ISA evaluation.
❷We formalize the analysis of commutation relations between arbitrary 2Q canonical gates that share one qubit. This offers a generalized commutativity-based optimization mechanism, moving beyond those tailored only for $\mathrm{CX}$ gates (liu2022not).
❸We conduct comprehensive experiments across a wide range of real-world benchmarks, hardware topologies, and representative ISAs, showing that Canopus consistently reduces routing overhead by 15%-35% compared to SOTA methods. Our results also yield holistic guidelines for the co-design of quantum programs, ISAs, and hardware.
❹We confirm that theoretically expressive ISAs exhibit superior performance to the conventional CX ISA, challenging the conclusions of prior works (kalloor2024quantum). We demonstrate some co-design guidelines for ISA-program-topology co-exploration.
❺Our case studies, including the real-machine QFT kernel execution and the end-to-end QEC circuit simulation, unequivocally showcase Canopus’ superiority in both near-term and fault-tolerant applications. For example, on the task of mapping QFT on 1D chain topology, Canopus finds the provably optimal routing scheme, surpassing the results previously reported as optimal in prior work (zhang2021time); and experiments on IBM’s QPUs suggest that Canopus leads to an average 2.1x fidelity improvement compared to Qiskit.
2. Background
2.1. Qubit mapping/routing
Figure 2. Mapping/routing to resolve topology constraints via $\mathrm{SWAP}$ insertion. With the initial mapping $\{q_{i}:Q_{i}\}$ (upper right), $g_{3}$ is not hardware compliant. Both $\mathrm{SWAP}_{q_{0},q_{1}}$ and $\mathrm{SWAP}_{q_{1},q_{2}}$ are sufficient to make $g_{3}$ executable.
Real quantum hardware typically have connectivity constraints, whereas algorithms often assume arbitrary interactions. To execute quantum circuits on topology-constrained hardware, logical qubits must first be mapped to physical qubit positions. This is called the initial mapping. In most cases, even an optimal initial mapping cannot guarantee all logical 2Q gates are mapped on physically connected qubit pairs. The common solution is to dynamically change logical-to-physical qubit mappings by inserting $\mathrm{SWAP}$ gates, as a $\mathrm{SWAP}$ gate exchanges state subspaces of two operand qubits, such that non-adjacent logical qubit states can be moved next to each other. Therefore, the qubit placement and routing compilation stage takes a logical circuit and hardware coupling graph as the input and outputs a transformed circuit within which each 2Q gate, with respect to a qubit mapping, is hardware compliant. An example is depicted in Figure 2.
2.2. Canonical description of 2Q gates
Figure 3. Geometric illustration of canonical gates confined to the Weyl chamber. For visualization convenience, herein the Weyl chamber is confined to $\left\{\frac{\pi}{4}\geq x\geq y\geq z\geq 0\right\}\cup\left\{\frac{\pi}{4}\geq\frac{\pi}{2}-x\geq y\geq z\geq 0\right\}$, equivalent to the canonical coefficient convention $\left\{(a,b,c)\,|\,\frac{1}{2}\geq a\geq b\geq|c|\right\}$ throughout this paper.
Any 2Q gate can be represented by a 4x4 matrix in $\mathbf{SU}(4)$, up to a global phase, with its canonical form defined as:
Definition 0 (Canonical gate).
Any 2Q gate $U\in\mathbf{SU}(4)$ can be expressed by the composition of its unique canonical form
(1) $\mathrm{Can}(a,b,c):=e^{-i\frac{\pi}{2}(a\,XX+b\,YY+c\,ZZ)},\,\frac{1}{2}\geq a\geq b\geq\lvert c\rvert$
sandwiched by local 1Q gates such that we say $U$ is locally equivalent to the canonical form $\mathrm{Can}(a,b,c)$.
The canonical coefficients $(a,b,c)$ are confined to a tetrahedron known as the Weyl chamber, which provides a geometric representation of all local equivalence classes of 2Q gates (zhang2003geometric). Figure 3 visualizes some common 2Q gates. E.g.,
  • $\mathrm{CX}$, $\mathrm{CZ}$, and $\mathrm{CR}$ are all equivalent to $\mathrm{Can}(\frac{1}{2},0,0)$.
  • $\mathrm{CX}$ family: $\mathrm{XX}(\theta)\sim\mathrm{YY}(\theta)\sim\mathrm{ZZ}(\theta)\sim\mathrm{Can}(\frac{\theta}{\pi},0,0)$.
  • Param-$\mathrm{SWAP}$ family: $\mathrm{pSWAP}(\theta)\sim\mathrm{Can}(\frac{1}{2},\frac{1}{2},\frac{1}{2}-\frac{\theta}{\pi})$.
In practice, the canonical form is acquired by KAK decomposition (tucci2005introduction) and has been widely used (bullock2003arbitrary; chen2024one).
2.3. Gate realization cost on hardware
The transformed circuits via qubit routing will be ultimately converted into basis gates for execution on hardware. Basis gates refer to those natively implemented and calibrated on physical platforms. Typical native gates in superconducting are $\mathrm{CR}$ (rigetti2010fully), $\mathrm{CZ}$, and $\mathrm{iSWAP}$ gates (krantz2019quantum). The realization cost of basis gates involves multiple aspects, including the benchmarked fidelity, gate duration, calibration efficiency, etc. For example, gates with shorter duration are more likely to achieve high fidelity, as qubit decoherence dominates the noise source; although some gate schemes can now implement more basis gates (chen2025efficient; nguyen2024programmable), those with simpler pulse control are more likely to be calibrated with high precision, such as the $\mathrm{iSWAP}$ family on flux-tunable transmons.
2Q gates that are not natively implemented must be synthesized by native gates. Their realization cost is determined by what basis gates are required for synthesis. For example, any 2Q gate can be minimally synthesized by 3 $\mathrm{CX}$ gates, except for $\mathrm{Can}(a,b,0)$ for which the required $\mathrm{CX}$ count is 2. Conventionally, $\mathrm{SWAP}$ is regarded as 3 times that of $\mathrm{CX}$ realization cost, while it can also be synthesized by “1 $\mathrm{CX}$ + 1 $\mathrm{iSWAP}$” or “3 $\mathrm{\sqrt{iSWAP}}$” gates. The monodromy polytope theory was proposed to determine the optimal synthesis cost for any 2Q gate given a specific set of basis gates through analysis of local invariants of canonical gates (peterson2020fixed). By this method, the set of gates realizable by a specified number of 2Q gates from the basis set, with arbitrary 1Q gates, corresponds to a polytope within the Weyl chamber. For instance, the polytope for 2 $\mathrm{\sqrt{iSWAP}}$ gates is a tetrahedron confined to $\left\{1/2\geq a\geq b+\lvert c\rvert\right\}$ (huang2023quantum).
3. Motivation
Limitations of conventional qubit routing models.
Conventional qubit routing models are ill-equipped to exploit the versatility of modern quantum hardware. First, whether optimizing for gate count or circuit depth, they typically assume that a $\mathrm{SWAP}$ costs 3 $\mathrm{CX}$ gates according to the textbook decomposition. This assumption is divorced from hardware reality. For example, a combination of $\mathrm{CX}$ and $\mathrm{iSWAP}$ is sufficient to realize a $\mathrm{SWAP}$ while both $\mathrm{CZ}$ (locally equivalent to $\mathrm{CX}$) and $\mathrm{iSWAP}$ are natively supported on mainstream superconducting platforms like Google’s Sycamore (arute2019quantum). Such platforms can even directly implement a high-fidelity $\mathrm{SWAP}$ gate, with the pulse duration $1.5$ times that of $\mathrm{CZ}$ (chen2025efficient). Thus, $\mathrm{SWAP}$ is not as costly as what these previous works on routing assume. Second, while prior works like liu2022not do assume the cost of a $\mathrm{SWAP}$ is context-dependent, their analysis remains confined to the $\mathrm{CX}$-based routing model. By relying on this overly simplistic model, conventional routers cannot accurately predict circuit execution costs and are blind to the significant optimization opportunities offered by richer, more diverse ISAs.
Co-optimization as the key to unlocking the superiority of advanced ISAs.
In response to the limitations of the $\mathrm{CX}$-only paradigm, a new generation of sophisticated quantum processors has emerged, featuring advanced ISAs with more powerful basis gates. Notable examples include the $\mathrm{\sqrt{iSWAP}}$ gate proposed by huang2023quantum, the continuous $\mathrm{ZZ}(\theta)$ (equivalent to $\mathrm{XX}(\theta)$, $\mathrm{ZX}(\theta)$, $\mathrm{MS}(0,0,\theta/2)$) jointly adopted by vendors (ibmFractionalGates; quantinuumArbitraryAngleGates; ionqPartialGates), and selected fractional or heterogeneous basis gates (mckinney2024mirage). Despite their theoretical promise for greater synthesis power and noise resilience, these advanced ISAs have largely remained in the proof-of-concept stage, with no systematic framework to harness their full potential in real-world quantum applications.
Prior efforts have been narrowly focused on local 2Q or multi-qubit synthesis tasks (huang2023quantum; tang2024quantum) or brute-force numerical optimizations (davis2019heuristics; younis2021qfast). Such rebase passes are tailored to a specific quantum ISA and fail to provide clear benefits when applied to advanced ISAs in realistic workloads (kalloor2024quantum). This has led to a critical question lingering in the community: “Are these more expressive, noise-resilient ISAs actually better?” Recently there have been attempts to harness the properties of advanced ISAs, although through manual and unsophisticated design, such as the $\mathrm{\sqrt{iSWAP}}$-based routing-synthesis optimization (mckinney2024mirage) and the $\mathrm{CX}$-$\mathrm{iSWAP}$ based routing for defect mitigation (zhou2024halma). In our work, we highlight that collaborative compiler optimization, especially at the stage following logical-level circuit optimization and followed by the final ISA rebase pass, is a key to fully exploit capabilities of those powerful ISAs: First, high-level algorithms are expressed in the $\mathrm{CX}$ representation, which then undergo template-based and peephole optimizations that are highly sophisticated and tailored for $\mathrm{CX}$-based circuit patterns (e.g., commutativity, Clifford equivalence); Second, the gap between the naïve qubit routing model and backend ISA properties apparently leaves a large untapped co-optimization space for both qubit routing and ISA rebase. Thus we are committed to validating this point through a systematic ISA-aware routing framework.
The “Tower of Babel dilemma” for utilizing diverse ISAs.
The proliferation of diverse quantum ISAs—from monolithic to complex, heterogeneous basis gate sets supported by various physical platforms—has created a “Tower of Babel dilemma” in the architecture and systems community. Developing bespoke compiler optimizations for each unique hardware backend is unsustainable, leading to the same software fragmentation that we have encountered in classical computing. Consequently, it is important to seek a unified approach that can effectively handle various platform-specific abstractions resembling the LLVM compiler (Lattner2004). The recently proposed monodromy polytope theory (peterson2020fixed), for example, provides the the method for optimal analysis of ISA synthesis capabilities. Specifically regarding the circuit-level compiler optimization, the monodromy polytope with canonical 2Q gate representations offers a unified approach to evaluating circuit cost and modeling routing-synthesis co-optimization. Building on this, Canopus offers an elegant and unified solution to the Tower of Babel dilemma at the compiler level.
Coherent cross-ISA, topology, and program pattern co-exploration.
Ultimately, the goal of quantum computing systems is not just to optimize software for existing hardware, but to co-design the entire stack—from algorithms to architecture—to build the most efficient system possible. This requires a holistic exploration of a vast and complex design space, asking critical questions like: Which ISA is best suited for a given class of applications (e.g., quantum error correction vs. quantum simulation)? How does the choice of qubit topology interact with the ISA to affect performance? Answering these questions is currently an ad-hoc, labor-intensive process, hindering systematic progress. Therefore, our work aims to provide the missing piece: a unified and automated framework for this co-exploration. By integrating qubit routing with a formal, ISA-aware synthesis cost model, Canopus can systematically evaluate the performance of different program patterns across heterogeneous ISAs and diverse hardware topologies. This enables researchers and hardware designers to make informed, data-driven decisions, accelerating the discovery of optimal co-design points and paving a systematic path toward robust, fault-tolerant quantum computer systems.
4. Canopus framework
4.1. Overview
Figure 4. Overview of the Canopus framework.
The overall routing procedure of Canopus is illustrated in Figure 4. The input circuit is rebased to $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$ before being fed to the routing pass. All processes operate on the directed acyclic graph (DAG) representation of the circuit. Canopus integrates the ISA-specific synthesis cost model into its $\mathrm{SWAP}$ search process and determines the best $\mathrm{SWAP}$ at each route step. The routing cost is efficiently computed by a formal analysis of 2Q canonical forms, without the ISA rebase step. Thus, the output is still a circuit DAG represented in $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$ with inserted $\mathrm{SWAP}$ gates.
Notably, Canopus inherits the basic concepts and data structures introduced in Sabre (li2019tackling), which is one of the industrial-standard qubit mapping algorithms. Given the input circuit DAG, Canopus first attempts to map 2Q gates layer by layer via extracting the front layer $F$, peeling executable gates and searching $\mathrm{SWAP}$ gates to minimize the routing cost according to a heuristic cost function. On the backbone of Sabre, we further introduce several key data structures such as the last mapped layer $L$ and the wire duration records $D$, to support the efficient implementation of ISA-aware routing and reducing both gate count and depth related routing overhead.
4.2. 2Q synthesis cost modeling
Figure 5. Synthesis coverage for $\bigl\{\mathrm{\sqrt{iSWAP}},\mathrm{ECP}\bigr\}$ gate set. The trivial points ($\mathrm{\sqrt{iSWAP}}$ and $\mathrm{ECP}$ themselves) are not shown in this figure. 2Q overage regions correspond to those that require (a) 2 $\mathrm{\sqrt{iSWAP}}$ gates or 2 $\mathrm{ECP}$ gates; (b) 1 $\mathrm{\sqrt{iSWAP}}$ + 1 $\mathrm{ECP}$; (c) 3 gates (3 $\mathrm{\sqrt{iSWAP}}$, 3 $\mathrm{ECP}$, 2 $\mathrm{\sqrt{iSWAP}}$ + 1 $\mathrm{ECP}$, etc.) from this gate set for synthesis, respectively.
As introduced in Section 2.3, given any basis gate set, the synthesis cost of a target 2Q gate can be exactly computed through the monodromy polytopes (peterson2020fixed). This cost (which basis gates are sufficient for synthesis) only depends on the canonical coefficients of the target gate. For example, Figure 5 illustrates various polytopes for the gate set $\bigl\{\mathrm{\sqrt{iSWAP}},\,\mathrm{ECP}\bigr\}$. In practice, the costs of each basis gate are pre-defined, thus the whole polytope helps decide the optimal synthesis scheme we should prioritize with the minimal circuit cost. For example, if $\mathrm{\sqrt{iSWAP}}$ and $\mathrm{ECP}$ have the same unit cost, the $\mathrm{SWAP}\sim\mathrm{Can}\bigl(\frac{1}{2},\frac{1}{2},\frac{1}{2}\bigr)$ gate realization will prioritize the “1 $\mathrm{\sqrt{iSWAP}}$ + 1 $\mathrm{ECP}$” combination; if $\mathrm{ECP}$ cost is set to be more than twice that of $\mathrm{\sqrt{iSWAP}}$, the $\mathrm{SWAP}$ realization prioritizes the “3 $\mathrm{\sqrt{iSWAP}}$” pattern.
4.3. Routing in canonical form
(a) ISA-aware $\mathrm{SWAP}$ insertion cost.
(b) $\mathrm{SWAP}$ insertion patterns with different gate count and depth costs.
Figure 6. Qubit routing with the canonical 2Q gate representation.
Our ISA-aware routing primarily leverages the mechanism that some inserted $\mathrm{SWAP}$ gates can “piggyback” a preceding 2Q gate with the same qubit pair acted on and thus result in lower (even negative) routing overhead than what naïve $\mathrm{SWAP}$ synthesis cost may imply. Based on the ISA-specific synthesis cost model, Canopus utilizes a holistic heuristic cost function that considers various requirements of qubit routing in a unified, quantitative approach.
As Figure 6(a) shows, every 2Q block within the circuit is consolidated into a single unitary gate and rebased to $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$. During qubit routing, when a $\mathrm{SWAP}$ insertion follows a 2Q gate $U$ acting on the same qubit pair, the actual insertion cost is reflected by the minimal set of basis gates sufficient for synthesizing the composite block $\mathrm{SWAP}\cdot U$, rather than by the cost of the $\mathrm{SWAP}$ gate itself ($c_{\mathrm{swap}}$). Specifically, Canopus tracks the set of 2Q canonical gates that have no succeeding 2Q gates within the DAG constructed by already routed gates, as the “last mapped layer” $L$. We define the cost component $c_{g}=\textsc{cost}(\mathrm{SWAP}\cdot U)-\textsc{cost}(U)$ if an inserted $\mathrm{SWAP}$ can be “absorbed” by $L$, while $c_{g}=c_{\mathrm{swap}}$ if it cannot. We further consider the cost for the circuit depth increment $\Delta_{\mathrm{depth}}$ during qubit routing, which is evaluated by tracking depth costs (durations) on every qubit wire, recorded as $D$. Figure 6(b) illustrates that different $\mathrm{SWAP}$ insertion patterns may lead to different gate count and depth costs, and thus should be comprehensively considered. Notably, the circuit depth (duration) is quantified with given basis gate costs, i.e., length of the weighted critical path on the circuit DAG. The detailed heuristic cost function in Canopus is defined as:
$H=\,\,$ $w_{g}\,c_{g}+w_{d}\,\Delta_{\mathrm{depth}}$
(2) $+(\Delta_{\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{F}}+k_{E}\,\Delta_{\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{E}})\,c_{\mathrm{swap}},$
where $w_{g}$ and $w_{d}$ are the weight factors for the gate count and depth cost components, respectively. The last term reflects the original heuristic cost $H_{\textsc{Sabre}}=\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{F}+k_{E}\,\mathrm{Avg}\{\mathrm{dist}[i,j]\}_{E}$ in Sabre (li2019tackling), while used in a unified approach accompanied with the other cost components in Canopus. Specifically, $H_{\textsc{Sabre}}$ refers to the average distance between physical qubits mapped to demanded logical interactions in the front layer $F$ and the extended set $E$ (lookahead mechanism). Here we use the “differential” average distance after inserting a $\mathrm{SWAP}$ candidate, then multiply it by the $c_{\mathrm{swap}}$ to reflect on average how an inserted $\mathrm{SWAP}$ can reduce the distance between each physical qubit pair, with the negative distance-cost increment quantified in the basis. Furthermore, the decay factor in Sabre is not needed anymore, as the $c_{g}$ and $\Delta_{\mathrm{depth}}$ guide more accurate co-optimization. Therefore, each cost component in Equation 2 refers to “cost increment”, composing a unified heuristic cost function to reconcile the multifaced routing costs.
4.4. Enhanced optimization via commutation
Figure 7. Canonical gate representation enables easily capturing commutative relations within real-world circuits. Commutativity within $\mathrm{CX}$ chain can be formally captured without tracking either’s control and target qubit positions. In the bottom subfigure $C$ indicates the 1Q Clifford gate.
Previous works have observed that employing the commutativity between $\mathrm{CX}$ gates exposes more optimization opportunities for $\mathrm{SWAP}$ insertion (liu2022not). However, the commutation pattern they exploit is limited to a pair of $\mathrm{CX}$ gates, where they either act on the same control qubit or target qubit. In our findings, the general 2Q gate commutativity can be captured through the canonical form:
Theorem 1 (Canonical gate commutation).
Let $\mathrm{Can}(a,b,c)_{q_{0},q_{1}}$ and $\mathrm{Can}(a^{\prime},b^{\prime},c^{\prime})_{q_{1},q_{2}}$ denote canonical gates acting on qubits ($q_{0},q_{1}$) and ($q_{1},q_{2}$) respectively, with an overlapping qubit $q_{1}$. They are commutative if and only if
(3) $b=b^{\prime}=c=c^{\prime}=0,$
that is, when both consist solely of $\mathrm{XX}$ rotations.
Detailed proof is provided in Appendix B. Through this formalized commutativity determination, the ordinary $\mathrm{CX}$ commutation pattern can be captured without tracking the control and target qubit positions, as shown in Figure 7.
4.5. Scalability and implementation
The overall algorithm framework to implement Canopus resembles Canopus. To efficiently implement the sophisticated $\mathrm{SWAP}$ insertion mechanism in Canopus, we develop corresponding core algorithms. Algorithm 1 specifies how the demanded data structures—the last mapped layer $L$, commutative canonical gate pairs $C$ within $L$, wire duration record DD—will be updated when adding an executable 2Q gate to the routed circuit DAG. Algorithm 2 shows how the wire durations $D$ should be correctly updated when encountering a $\mathrm{SWAP}$ insertion that satisfies the canonical gate commutativity optimization opportunity. It is also essential to evaluate the total circuit cost after mapping. Notably, all the computation processes within these algorithms are based on conditional control and operations on hashed data structures, thus with $\mathcal{O}(1)$ time complexity introduced. To quantify the synthesis cost computation of a given canonical gate, the computation is to determine which convex polytope its canonical coordinate is located in, which is a highly efficient computation process with linear time complexity. Canopus will also cache canonical gate costs it has acquired to avoid repetitive computation. Consequently, the overall scalability of Canopus is on par with that of Sabre, ensuring its practical applicability to large-scale circuits.
1
2
Input :  $G^{\prime}$ (Routed DAG), $\pi$ (current logic-to-physical mapping), $L$ (last mapped layer), $D$ (wire durations for each qubit), $C$ (commutative pairs within $L$)
Output : Updated $G^{\prime}$, $L$, $D$, $C$
3
/* $g$: resolved logical gate;  $g^{\prime}$: routed gate */
$g^{\prime}\leftarrow G^{\prime}.\textsc{pushBack}(g,\pi[g.q_{0}],\pi[g.q_{1}])$;  // $g^{\prime}.q_{i}=\pi[g.q_{i}]$
4
[…]
6 $D[g^{\prime}.q_{0}]\leftarrow d$; $D[g^{\prime}.q_{1}]\leftarrow d$;
7
8for $\mathrm{pred}\in G^{\prime}.\textsc{predecessors}(g^{\prime})$ do
9 if $\textsc{is2QGate}(\mathrm{pred})$ then
10    if $\textsc{isCommutativeCanonicalPair}(g^{\prime},\,\text{pred})$ then
11       $C[(\mathrm{pred}.q_{0},\,\text{pred}.q_{1})]\leftarrow(g^{\prime}.q_{0},\,g^{\prime}.q_{1})$;
12
13    else
14       $L.\textsc{pop}((\mathrm{pred}.q_{0},\,\mathrm{pred}.q_{1}),\,\textsc{None})$;
15       $C.\textsc{pop}((\mathrm{pred}.q_{0},\,\mathrm{pred}.q_{1}),\,\textsc{None})$;
16
17
18 else
    /* $\mathrm{pred\_pred}$ must be None or a 2Q gate */
19    $\mathrm{pred\_pred}\leftarrow\textsc{next}(G^{\prime}.\textsc{predecessors}(\mathrm{pred}))$;
20    if $\mathrm{pred\_pred}\neq\textsc{None}$ then
21       $L.\textsc{pop}((\mathrm{pred\_pred}.q_{0},\,\mathrm{pred\_pred}.q_{1}),\,\textsc{None})$;
22       $C.\textsc{pop}((\mathrm{pred\_pred}.q_{0},\,\mathrm{pred\_pred}.q_{1}),\,\textsc{None})$;
23
24
25
[…]
27
Algorithm 1 Update $L$ when adding a new 2Q gate
1
2
Input :  swap (encountered SWAP gate), can (canonical gate within $L$ on the same qubits as swap), $D$, $C$
Output : Updated $D$
3
4if $(\texttt{swap}.q_{0},\texttt{swap}.q_{1})\in C$ then
5 $q^{\prime}_{0},\,q^{\prime}_{1}\leftarrow C[(\texttt{swap}.q_{0},\,\texttt{swap}.q_{1})]$;
 /* Adjust $D$ by finding matched qubits $q_{i}\in\{\texttt{swap}.q_{0},\texttt{swap}.q_{1}\}$ and $q^{\prime}_{j}\in\{q^{\prime}_{0},q^{\prime}_{1}\}$ */
6 $D[q_{i}]\leftarrow D[q^{\prime}_{j}]+\textsc{synthCost}(\text{can})$;
[…]
8
9
[…]
11 $D[\texttt{swap}.q_{0}]\leftarrow d$; $D[\texttt{swap}.q_{1}]\leftarrow d$;
12
Algorithm 2 Update $D$ when adding a $\mathrm{SWAP}$ gate
5. Case Studies
We validate the practical advantages of Canopus through two realistic case studies: the real-machine compilation and execution of Quantum Fourier Transform (QFT) circuits on IBM’s QPU ibm_torino, and the end-to-end simulation of compiled Quantum Low-Density Parity-Check (QLDPC) stabilizer circuits to assess its impact on the logical error rate.
5.1. QFT kernel
(a) Mapping qft_6 by TOQM.
(b) Mapping qft_6 by Canopus.
Figure 8. Mapping/routing comparison for the QFT kernel. For convenient visualization, only $\mathrm{CPhase}$ and $\mathrm{SWAP}$ gates are shown. (a) TOQM generates a sub-optimal mapping scheme, with 2Q depth of 10. (b) Canopus generates the optimal scheme in a perfect butterfly structure, with 2Q depth of 9.
Table 1. Qubit routing comparison for the QFT kernel.
QFT kernel qft_6 qft_12
Topology Method #Can Depth2Q #Can Depth2Q
1D Chain Optimal 15 9 66 21
TOQM 16 10 67 22
Canopus 15 9 66 21
2D Square TOQM 21 13 100 39
Canopus 15 9 75 (±10%) 33 (±10%)
QFT is a fundamental subroutine in many promising quantum algorithms, such as Shor’s algorithm (shor1994algorithms), quantum phase estimation (kitaev1995quantum), etc. Given the extensive research on dedicated QFT compilers (zhang2021time; jin2024optimizing; maslov2007linear), we select the state-of-the-art compiler TOQM, which specializes in QFT optimization, as our primary baseline.
A key finding is that Canopus always achieves the optimal QFT routing scheme on the 1D chain topology, while TOQM does not. Provably, the minimal number of $\mathrm{SWAP}$ insertions to route an $n$-qubit QFT is $\frac{n(n-1)}{2}-2$, that is, 2 fewer than the original $\mathrm{CPhase}$ count. This results in a perfect, symmetric butterfly circuit structure, as exemplified in Figure 8(b), with minimal #Can and 2Q circuit depth. Notably, this result is indeed optimal, surpassing the manually designed scheme previously reported as optimal by maslov2007linear where 2 more $\mathrm{SWAP}$ gates are required. This optimal scheme is irrespective of the target ISA. In contrast, our experiments show that TOQM  despite claiming to realize the scheme from (maslov2007linear), fails to reproduce it and consistently yields inferior results to Canopus, as illustrated in Figure 8.
We compare compilation performance for both $6$- and $12$-qubit QFT kernels on both 1D chain and 2D square topologies, with results summarized in Table 1. On the 1D chain, Canopus always produces the theoretically optimal routing result, while TOQM does not. For the small-scale qft_6 kernel on the 2D square, Canopus also achieves the optimal routing, superior to TOQM in both #Can and 2Q depth. For the large-scale qft_12 kernel, Canopus consistently outperforms TOQM in both metrics.
Figure 9. QFT kernel fidelity comparison benchmarked on IBM® Quantum Platform (ibm_torino). ibm_torino is the Heron-series QPU with native gate set $\bigl\{\mathrm{CZ},\,\sqrt{\mathrm{X}},\,\mathrm{Z}(\theta)\bigr\}$.
To further validate these results, we performed real-machine experiments on IBM’s ibm_torino QPU. We compiled QFT circuits of sizes $n\in\{6,8,10,12\}$ for a 1D chain topology using both Canopus and the default Qiskit compiler. Although ibm_torino has a heavy-hex topology, it contains linear chains of sufficient size for these benchmarks. Fidelity was measured using the Hellinger fidelity between the experimental and ideal output distributions, with the number of shots set to $\textsc{max}\{4096,2^{n}\times 10\}$. A layer of Hadamard gates is appended to each circuit execution so that the ideal final state will be $\lvert 0\rangle^{\otimes n}$. In Figure 9, circuits compiled with Canopus achieve, on average, a 52.9% reduction in $\mathrm{CZ}$ gate count, a 66.4% reduction in 2Q-gate depth, and a 2.1x improvement in fidelity. These results unequivocally demonstrate the practical advantages of Canopus for QFT kernel compilation.
5.2. QLDPC stabilizer circuit
Figure 10. Logical error rate of QLDPC stabilizer circuits compiled for square (top) and heavy-hex (bottom) topologies. The y-axis shows the relative logical error rate normalized by the error rate of an ideal baseline that assumes all-to-all connectivity.
QLDPC codes are rapidly moving from a topic of theoretical interest to a cornerstone of experimental fault-tolerant quantum computing (FTQC) research, mainly because of their superior encoding efficiency (bravyi2024high; breuckmann2021quantum; Dennis_2002). However, due to their frequent long-range interactions (breuckmann2021quantum; panteleev2021degenerate), realizing QLDPC codes on superconducting processors with fixed, local connectivity is still hampered by significant routing overheads (wang2025demonstration).
Therefore, we attempt to compile the circuits automatically under two ISAs: (1) CX ISA with $\mathrm{CX}$ as the 2Q basis gate; (2) Stab ISA with both $\mathrm{CX}$ and $\mathrm{iSWAP}$ as basis gates, assumed to have an identical cost. Particularly, the Stab ISA aligns with practical hardware realities, e.g., both $\mathrm{CZ}$ and $\mathrm{iSWAP}$ can be natively supported by mainstream superconducting platforms (krantz2019quantum; wei2024native; arute2019quantum; nguyen2024programmable). We further build an end-to-end evaluation pipeline with QLDPC code circuit examples from (wang2025demonstration; panteleev2021degenerate). We simulate the routed circuits using stim (Gidney2021StimAF) under the circuit-level noise model (Dennis_2002; Acharya2022SuppressingQE). Finally, all syndromes are decoded using the BPOSD decoder (panteleev2021degenerate; roffe2020decoding; hillmann2024localized) to determine the logical error rate.
As shown in Figure 10, Canopus consistently achieves lower logical error rates than Sabre, as the ISA-aware approach of Canopus results in compiled circuits with less $\mathrm{CX}/\mathrm{iSWAP}$ gate count and circuit depth. Under the CX, Canopus yields an average logical error suppression of 49.4% on the square topology and 11.4% on the heavy-hex topology compared to Sabre. The advantage becomes even more pronounced with the Stab, where Canopus achieves a 52.6% (square) and 29.3% (heavy-hex) error suppression. These results highlight two key findings: first, the ISA-aware mechanism in Canopus is highly effective for compiling QEC circuits, and second, the dedicated use of a hybrid $\mathrm{CX}$-$\mathrm{iSWAP}$ gate set offers a significant practical advantage for QLDPC code demonstrations on superconducting hardware.
6. Evaluation
We further holistically evaluate Canopus compared to other leading methods, across representative ISAs and hardware topologies. The evaluation provides both cross-compiler and cross-ISA comparisons under the coherent settings for basis gate cost and routing overhead metric.
Table 2. Selected quantum ISAs.
ISA 2Q basis gates Description
CX $\left\{\text{CX}\right\}$ Conventional CX gate
ZZPhase $\bigl\{\mathrm{ZZ}_{\frac{\pi}{6}},\mathrm{ZZ}_{\frac{\pi}{4}},\mathrm{ZZ}_{\frac{\pi}{2}}\bigr\}$ Discrete CX-family gates, i.e., $\bigl\{\sqrt[3]{\mathrm{CX}},\sqrt{\mathrm{CX}},\mathrm{CX}\bigr\}$ (peterson2022optimal)
SQiSW $\bigl\{\mathrm{\sqrt{iSWAP}},\mathrm{iSWAP}\bigr\}$ Half evolution of $\mathrm{iSWAP}$ and $\mathrm{iSWAP}$ (huang2023quantum)
ZZPhase_ ZZPhase++ $\bigl\{\mathrm{pSWAP}_{\frac{\pi}{6},\,\frac{\pi}{4},\,\frac{\pi}{2}}\bigr\}$ ZZPhase ISA with the mirror gates
SQiSW_ SQiSW++ $\left\{\mathrm{ECP},\mathrm{CX}\right\}$ SQiSW ISA with the mirror gates (mckinney2024mirage)
Het ZZPhase++ SQiSW Heterogeneous CX-family and iSWAP-family gates
Table 3. Benchmarks information. These metrics are collected from TKet-optimized logical circuits with only $\mathrm{Can}$ and $\mathrm{U3}$ gates. Circuit cost ($C_{\mathrm{count}}$ and $C_{\mathrm{depth}}$) is calculated in CX ISA.
Program #Qubit #Can Depth2Q $C_{\mathrm{count}}$ $C_{\mathrm{depth}}$
bigadder (li2023qasmbench) 18 114 79 130.0 88.0
bv (li2023qasmbench) 19 18 18 18.0 18.0
ising (li2023qasmbench) 26 25 2 50.0 4.0
knn (li2023qasmbench) 25 72 50 84.0 62.0
multiplier (li2023qasmbench) 15 198 122 222.0 133.0
qec9xz (li2023qasmbench) 17 32 12 32.0 12.0
qft (quetschlich2023mqt) 18 153 33 306.0 66.0
qpeexact (quetschlich2023mqt) 16 127 43 260.0 86.0
qram (li2023qasmbench) 20 110 70 130.0 78.0
sat (li2023qasmbench) 11 210 182 252.0 204.0
swap_test (li2023qasmbench) 25 72 50 84.0 62.0
wstate (li2023qasmbench) 27 52 28 52.0 28.0
6.1. Experimental settings
6.1.1. ISAs and basis gate costs
We consider six different ISAs (including the conventional CX ISA) listed in Table 2. These cover a wide range of powerful basis gates from $\mathrm{CX}$-family and $\mathrm{iSWAP}$-family gates. Particularly, SQiSW (huang2023quantum) proves to a powerful ISA option and has been adopted by recent software projects (mckinney2024mirage; cirqSQiSWDecomposer). ZZPhase ISA containing three fractional $\mathrm{ZZ}(\theta)$ rotation gates (equivalently, $\left\{\sqrt[3]{\mathrm{CX}},\sqrt{\mathrm{CX}},\mathrm{CX}\right\}$) is adopted by Qiskit’s latest synthesis functionalities (peterson2022optimal; qiskitXXDecomposer). For ZZPhase and SQiSW, we also consider the mirror-enhanced version by incorporating the mirrored basis gates (cross2019validating; mckinney2024mirage) into the ISAs. We also include the Het ISA that is the composition of ZZPhase and SQiSW.
To conduct a coherent cross-ISA performance comparison, we use a consistent basis gate cost setting:
(6) $\left\{\begin{array}[]{c}\mathrm{CX}:1,\,\mathrm{ZZ}(\pi/t):2/t,\,\mathrm{\sqrt{iSWAP}}:0.75,\\ \mathrm{iSWAP}:1.5,\,\mathrm{ECP}:1.25,\,\mathrm{pSWAP}(\pi/t):2-1/t\end{array}\right\},$
where $\mathrm{CX}$ gate is the unit cost. Such a setting ensures the continuity of gate costs along the critical edges in the Weyl chamber. For example, $\mathrm{pSWAP}(\pi/2)$ is equivalent to $\mathrm{iSWAP}$ and they have the same cost of $1.5$. With a specific gate family, basis gates with larger canonical coefficients usually requires proportionally longer interaction time on physical qubits, which was reflected in the cost setting. Note that this setting is a comprehensive consideration for current gate schemes and hardware-implemented gate fidelities in superconducting (chen2025efficient; arute2019quantum; wei2024native; nguyen2024programmable; acharya2024quantum). It is neither limited to a specific gate scheme nor a specific hardware platform.
6.1.2. Metrics
With the consistent basis gate cost settings above, we can evaluate cross-ISA circuit cost comparison, in terms of both gate count ($C_{\mathrm{count}}$) and circuit depth ($C_{\mathrm{depth}}$). Specifically, $C_{\mathrm{count}}$ refers to the sum of all 2Q gate costs according to the basis gate setting in Equation 6. $C_{\mathrm{depth}}$ refers to the length of the cost-weighted critical path within the circuit DAG. $C_{\mathrm{count}}$ and $C_{\mathrm{depth}}$ are naturally the generalized metrics for 2Q gate count and circuit depth. To quantify the routing effects across ISAs and topologies, we define the routing overhead as the ratio of routed circuit cost to the pre-routed circuit cost, for which the pre-routed logical-level circuit cost is uniformly computed in the CX ISA.
6.1.3. Benchmarks
We select a set of medium-size benchmarks from QASMBench (li2023qasmbench) and MQTBench (quetschlich2023mqt) spanning various categories of quantum programs. These benchmarks first go through logical-level optimization by TKet and are rebased to $\left\{\mathrm{Can},\,\mathrm{U3}\right\}$ as the input of the evaluated compilers. Information for benchmarks after logical optimization is summarized in Table 3.
6.1.4. Baselines
The leading methods Sabre, TOQM, and BQSKit are selected as the baselines. Both Sabre and Canopus are implemented in the Python-based Qiskit framework, i.e., we do not use the Rust-accelerated Sabre implementation in the latest Qiskit version, for fair runtime comparison. TOQM is the SOTA circuit depth driven qubit routing method (zhang2021time). We also select BQSKit as a baseline as it represents another different cross-ISA compilation paradigm (bqskit). Given a target gate set and coupling graph, BQSKit performs end-to-end compilation via numerical optimization, that is, finally the rebased circuit is generated.
Hyperparameters for Sabre and Canopus are of the same settings. Each performs 10 times layout procedure, within which 5-round bidirectional passes are proceeded and each pass performs 10 trials. The best result across all attempts is selected. TOQM can obtain the deterministic routing result in one go. Compiled circuits by BQSKit, although in terms of only the 2Q gates, is also random. Thus we perform 3 trials for each input case and report the best result.
6.2. Suppression of routing overhead
Table 4. Average (geometric-mean) routing overhead across different topologies and backend ISAs.
Routing overhead In terms of $C_{\mathrm{count}}$ In terms of $C_{\mathrm{depth}}$
Topo ISA Type sabre toqm bqskit canop sabre toqm bqskit canop
[…]
The comprehensive evaluation (on geometric-mean average) results are summarized in Table 4, across the six selected ISAs and three hardware topologies (1D chain, 2D heavy-hex, 2D square). While detailed results across the $3\times 6\times 12$ cases are listed in Table 5 and Table 6 in Appendix C. Canopus results in the lowest average (geometric-mean) routing overhead for each ISA-topology combination benchmarking, significantly surpassing other methods. Specifically, the routing overhead reduction achieved by Canopus is on average 16.06% (26.44%), 34.70% (21.25%), and 19.89% (20.72%) in terms of $C_{\mathrm{count}}$ ($C_{\mathrm{depth}}$), compared to Sabre, TOQM, and BQSKit, respectively. With Canopus, those more powerful ISAs, such as ZZPhase and SQiSW ISAs, actually exhibit significantly lower routing overhead than CX. While the baseline methods do not apparently indicate that, for which the circuit cost reduction stems from ISA rebase itself.
Notably, Sabre and BQSKit focus on gate count driven optimization, TOQM specializes in lowering circuit depth, while our Canopus involves both count and depth related optimization. Among them, TOQM leads to the worst count-related routing overhead. Even for the routing overhead in terms of depth, TOQM consistently underperforms Canopus, and it only outperforms BQSKit for 2D square topology. Besides, Canopus could manage the routing overhead in a ubiquitously low level, in terms of both count and depth costs; while every baseline fails to optimize some specific circuits. For example, TOQM/BQSKit cannot manage the routing overhead for qec9 circuit; BQSKit cannot for bv circuit, even though with more expressive ISAs.
6.3. Program-ISA-Topology co-exploration
Our evaluation also systematically explores how program patterns, ISA selection, and hardware topologies impact each other. We highlight some guidelines for optimal co-design particularly according to results achieved by Canopus (Table 4, Table 5, Table 6): ① Heavy-hex topology consistently leads to higher routing overhead across all these ISAs, compared to the seemingly more sparse 1D chain topology. This is mostly because most quantum algorithms are constructed in a subroutine-unrolling approach, naturally more friendly with chain topology. The QFT kernel detailed in Section 5.1 is a thorough good example. ② The combinatorial ISA involving both $\mathrm{CX}$-family and $\mathrm{iSWAP}$-family gates is much superior to either sole one. Specifically, either ZZPhase or SQiSW could lead to no more than 11% routing overhead reduction compared to CX, while Het results in more than 25% reduction than CX. This benefit is more significant for circuits largely containing $\mathrm{CX}/\mathrm{CZ}$ as 2Q blocks such as qec9. ③ Gate mirroring is another approach to designing powerful quantum ISAs, as both ZZPhase and SQiSW achieve comparable results to Het ISA. This is due to the same optimization opportunities brought by $\mathrm{SWAP}$ mirroring. ④ For Hamiltonian simulation programs like ising, ZZPhase is essential to improve execution performance. Therein multiple 2-local Pauli rotations equivalent to $\mathrm{XX}(\theta)$ are included. As a discrete fractional $XX(\theta)$ basis gate set, ZZPhase ISA is more suitable than other gate family selections in this case.
The real-machine experiment in Section 5.1 showcases how our method can help achieve superior compilation results and thus higher program fidelities for QFT kernels using the standard CX ISA ($\mathrm{CZ}$ on ibm_torino). However, extending this real-machine validation to alternative ISAs is currently challenging due to the scarcity of quantum processors equipped with well-calibrated heterogeneous gate sets. For instance, while IBM has proposed fractional gates, i.e., the continuous $\mathrm{ZZ}(\theta)$ gate set (ibmFractionalGates), their implementation details and calibration procedures are not publicly disclosed. To our knowledge, the $\mathrm{ZZ}(\theta)$ gates have the same duration as $\mathrm{CZ}$ on IBM’s Heron QPUs regardless of the rotation angle, and their error rates are consistently 1x-3x that of $\mathrm{CZ}$. This performance is far from the ideal assumptions of ZZPhase. Fortunately, a path forward is emerging with the recently proposed AshN gate scheme (chen2024one) that enables directly implementing any basis gates with the optimal gate durations. It is also experimentally evinced on transmon qubits by chen2025efficient, where multiple basis gates are calibrated in high fidelities that aligns with our cost model as well. This development may enable comprehensive, real-machine co-exploration of programs, ISAs, and hardware topologies in the near future.
6.4. Diverse-ISA compilation paradigms
Prior to this work, there are two major compilation paradigms targeting diverse ISAs: (1) Use the conventional compiler that operates entirely on the $\mathrm{CX}$-based circuit representation before ISA rebase. The final-stage rebase pass can usually be completed via optimal synthesis in efficient analytical or numerical computation (tucci2005introduction; peterson2022optimal; huang2023quantum; mckinney2025two). (2) Use brute-force approximate synthesis to perform structural search and numerical optimization to determine the synthesized circuit with minimal gate count (davis2019heuristics; kukliansky2023qfactor; patel2022quest). Sabre/TOQM and BQSKit are representative of these two paradigms, respectively. In our evaluation, BQSKit even underperforms the industrial-standard Sabre in most cases. Exceptionally, in terms of the circuit depth, BQSKit leads to better results than other baselines on sparse topologies (chain, heavy-hex), as its A*-based search for 2Q gate arrangement could exhibit advantages over long-range qubit routing, but this advantage does not hold for more connected topologies. Besides, the second numerical optimization based paradigm is of exponential computational complexity. For benchmarking the 216 medium-size cases in Table 5, the Rust-backend BQSKit requires on average 18 minutes to process each circuit with an Apple M3 Max CPU; in contrast, the Python-implemented Sabre requires only 17 seconds. Consequently, this second paradigm is ill-suited for compiling real-world programs, proving both ineffective and inefficient when targeting diverse ISAs (at least for discrete gate sets). Instead, although there is a gap between the conventional routing model and backend ISA properties, by means of routing-synthesis co-optimization, the first paradigm could be enhanced to bridge the gap between the routing model and backend ISA properties and thus provides a more viable path.
6.5. Runtime analysis
Figure 11. Compilation latency comparison.
In our field tests for the set of benchmarks above, Canopus consistently exhibits 1x-2x runtime latency of Sabre. This result aligns with the complexity analysis in Section 4.5. Herein we specifically demonstrate the end-to-end runtime analysis for larger-scale quantum circuits. We use random quantum volume (cross2019validating) circuits generated by Qiskit for scalability benchmarking, which represent the most complicated circuit structures with dense 2Q gate arrangement. Each canonical gate within the quantum volume circuit contains unique canonical parameters as each 2Q unitary is randomly generated, thus there is no cached synthesis cost calculation for performance improvement in one pass. We select quantum volume circuits with two different widths (number of qubits), 15 and 20. We vary the depth of these circuits (qv_15 and qv_20) from 50 to 200. The largest size of these qv circuits is up to thousands of 2Q gates. As Figure 11 shows, Canopus leads to an average 1.31x (±1%) latency compared to Sabre for each benchmarked circuit. Both compilers’ latency scales linearly with circuit depth and width. If we compare the curve slopes, Canopus leads to 1.32x (1.30x) latency scaling slope compared to that of Sabre in terms of depth for qv_15 (qv_20) circuits.
7. Related Work
Qubit mapping/routing is one of the most well-explored topics of quantum compiler research, as it shares similar methodologies with instruction scheduling (codina2001unified; hennessy1983postpass) and register allocation (chaitin1982register; poletto1999linear) in classical computing.
To perform scalable qubit routing, zulehner2018efficient introduces an A*-based algorithm to minimize $\mathrm{SWAP}$ gate overhead for concurrent $\mathrm{CX}$ gate layers. The approach partitions the circuit into layers and solves the mapping problem subsequently. li2019tackling also utilizes the circuit DAG layering thought and proposes a bidirectional routing procedure Sabre to find better initial mappings thus with lower $\mathrm{SWAP}$ insertion count. It also briefly discusses the trade-off between the inserted $\mathrm{SWAP}$ count and the circuit depth but does not prioritize optimizing circuit depth. Subsequent works have aimed to improve circuit depth and parallelism, either by using Sabre-like heuristics (lao2021timing; ddroute2025; zou2024lightsabre) or graph matching techniques (childs2019circuit). zhang2021time systematically investigates the depth-optimality of qubit mapping and proposes an A*-based method TOQM that reported superior performance over existing solver-based depth-driven approaches (tan2020optimal). However, holistic optimality of qubit routing is contingent on the specific ISA, device topology, and circuit cost model, and is rarely guaranteed by theoretical bounds. Indeed, our own evaluation reveals that TOQM does not always produce depth-optimal results compared to our heuristic, Canopus. For example, our case study in Section 5.1 demonstrates that the mapping scheme for the QFT kernel, purported to be optimal in their analysis, can be further improved.
With the recent development of advanced quantum ISAs such as superconducting fractional gates (ibmFractionalGates), ion-trapped partial entangling gates (ionqPartialGates; yale2025realization), and the AshN gates (chen2024one; chen2025efficient), some works have begun exploring how to efficiently utilize these ISAs to make compiler optimizations closer to hardware characteristics. mckinney2024mirage investigates the practical performance of SQiSW ISA proposed by huang2023quantum and the synthesis capability when incorporating the basis gates’ mirrors into the ISA. Their modified Sabre algorithm offers a preliminary attempt at the collaborative gate decomposition and qubit routing approach, while the optimization opportunities considered therein are limited and the algorithmic techniques are not sophisticated. BQSKit (bqskit) and the series of works behind it (davis2019heuristics; wu2020qgo; kukliansky2023qfactor; younis2021qfast) provide a toolkit to rebase arbitrary 2Q unitaries to specific ISAs through approximate synthesis (structural search and numerical optimization) which is not computationally efficient. Approximate synthesis by BQSKit does not ensure optimal schemes for two-qubit and multi-qubit circuit synthesis. In addition, due to the lack of native compilation strategies and a rational synthesis cost model, kalloor2024quantum claims that alternative ISAs are hardly comparable to CX when evaluating quantum hardware roofline by BQSKit. As for applicability of expanded ISAs to QEC, Google’s latest theoretical (mcewen2023relaxing) and experimental (eickbusch2024demonstrating) works demonstrate the CX-iSWAP combination ISA could help suppress the fault-tolerant threshold. zhou2024halma proposes a routing-based method enhanced by CX-iSWAP for overcoming ancilla defects among surface code blocks while preserving encoded logical information, but it relies on manual design and experience.
8. Conclusion
In our work, we introduce Canopus, the first unified, ISA-aware qubit routing framework designed to operate across diverse quantum hardware. By leveraging a canonical two-qubit gate representation and a formal cost model derived from monodromy polytope theory, Canopus achieves deep co-optimization of qubit routing and gate synthesis. This approach not only demonstrates the practical superiority of emerging quantum ISAs but also enables systematic co-exploration of how different ISAs, program patterns, and hardware topologies interact, providing a powerful new tool for quantum system design.
Appendix A Canonical gate and 2Q circuit synthesis
In this section we show the basic mathematical properties of the canonical form of 2Q unitary and then discuss the synthesis capability of some 2Q basis gates.
A.1. Canonical decomposition
$\mathbf{SU}(N)$ is a real manifold with dimension $N^{2}-1$, within which any element is a special unitary matrix with determinant equal to 1. Since the global phase does not affect quantum computation processes, it is sufficient to focus on the mathematical properties of special unitaries in the area of circuit synthesis. A generic 2Q gate, despite having 15 real parameters, can have its nonlocal behavior fully characterized by only 3 real parameters. This method, known as Canonical decomposition or KAK decomposition from Lie algebra theory, is widely adopted in quantum computing (zhang2003geometric; tucci2005introduction; bullock2003arbitrary; zulehner2019compiling). Specifically, for any $U\in\mathbf{SU}(4)$, there exists a unique $\vec{\eta}=(x,y,z)\in W\subseteq\mathbb{R}^{3}$, along with $V_{1},V_{2},V_{3},V_{4}\in\mathbf{SU}(2)$ and a global phase, such that
(7) $U=g\cdot(V_{1}\otimes V_{2})e^{-i\vec{\eta}\cdot\vec{\Sigma}}(V_{3}\otimes V_{4}),\,g\in\{1,i\}$
where $\vec{\Sigma}\equiv(XX,YY,ZZ)$ (tucci2005introduction). The set
(8) $W:=\left\{(x,y,z)\in\mathbb{R}^{3}\,|\,\frac{\pi}{4}\geq x\geq y\geq|z|,\,z\geq 0\text{ if }x=\frac{\pi}{4}\right\}$
is known as the Weyl chamber (zhang2003geometric), and $\vec{\eta}\in W$ is known as the Weyl coordinate of $U$. We also refer to a gate of the form
(9) $\mathrm{Can}(a,b,c):=e^{-i\frac{\pi}{2}(a\,XX+b\,YY+c\,ZZ)}=\begin{pmatrix}e^{-i\frac{c\pi}{2}}\cos{\frac{(a-b)\pi}{2}}&0&0&-ie^{-i\frac{c\pi}{2}}\sin{\frac{(a-b)\pi}{2}}\\ 0&e^{i\frac{c\pi}{2}}\cos{\frac{(a+b)\pi}{2}}&-ie^{i\frac{c\pi}{2}}\sin{\frac{(a+b)\pi}{2}}&0\\ 0&-ie^{i\frac{c\pi}{2}}\sin{\frac{(a+b)\pi}{2}}&e^{i\frac{c\pi}{2}}\cos{\frac{(a+b)\pi}{2}}&0\\ -ie^{-i\frac{c\pi}{2}}\sin{\frac{(a-b)\pi}{2}}&0&0&e^{-i\frac{c\pi}{2}}\cos{\frac{(a-b)\pi}{2}}\end{pmatrix}$
as a canonical gate. Two 2Q gates $U$ and $V$ are considered locally equivalent if they differ only by 1Q gates, meaning their canonical coefficients can be transformed into one another via the equivalence rules (crooks2020gates):
  1. (1)
    $(a,b,c)\sim(b,a,c)$ or $(a,b,c)\sim(c,b,a)$, i.e., any permutation of the coefficients;
  2. (2)
    $(a,b,c)\sim(-a,-b,c)$;
  3. (3)
    $(a,b,c)\sim(a-1,b,c)$;
  4. (4)
    $(1/2,b,c)\sim(1/2,b,-c)$.
Note that we align the conventional that canonical coefficient $(a,b,c)$ differs from Weyl coordinate $(x,y,z)$ by a $\frac{\pi}{2}$ factor. Unless otherwise specified, the canonical coefficients of gates in quantum ISAs and circuits are confined to $\frac{1}{2}\geq a\geq b\geq\lvert c\rvert$. While for the Weyl chamber visualization by means of weylchamber (weylchamber), we assume the Weyl coordinates are confined to $\left\{\frac{\pi}{4}\geq x\geq y\geq z\geq 0\right\}\cup\left\{\frac{\pi}{4}\geq\frac{\pi}{2}-x\geq y\geq z\geq 0\right\}$, as illustrated by Figure 3. Conversion of Weyl coordinates for different conventions is simple according to the equivalence rules above.
A.2. Quantum ISA and the synthesis capability
Figure 12. Coverage set for CX ISA.
Figure 13. Coverage set for SQiSW ISA.
Figure 14. Coverage set for SQiSW_ ISA.
Figure 15. Coverage set for ZZPhase ISA.
Figure 16. Coverage set for ZZPhase_ ISA.
Figure 17. Coverage set for Het ISA.
A quantum ISA typically includes qubit initialization, a universal gate set, and measurement. It serves as an interface between software and hardware by mapping high-level semantics of quantum programs to low-level native quantum operations or pulse sequences on hardware. The universal gate set, especially specified by its 2Q basis gates, is the key component of a quantum ISA that dominates its hardware-implementation accuracy and cost, as well as software-expressivity sufficiency.
$\mathrm{CX}$ or $\mathrm{CNOT}$ is the most popular basis gate provides by hardware vendors and considered by various quantum compiler optimization methods. The superconducting Cross-Resonance gate (rigetti2010fully) and ion-trapped Mølmer-Sørensen gate (bruzewicz2019trapped) are both $\mathrm{CX}$-equivalent gates with the same canonical form $\mathrm{Can}\bigl(\frac{1}{2},0,0\bigr)$. In the superconducting platforms with $XY$-coupled Hamiltonian like Google’s Sycamore (arute2019quantum), $\mathrm{iSWAP}\sim\mathrm{Can}\bigl(\frac{1}{2},\frac{1}{2},0\bigr)$ is another representative native 2Q basis gate and could be less sensitive to leakage error than the native $\mathrm{CZ}$ gate. Recent experimental advances demonstrate that more basis gates could be implemented natively and calibrated in high precision (chen2025efficient; wei2024native; yale2025realization). Particularly, some basis gates like $\mathrm{\sqrt{iSWAP}}\sim\mathrm{Can}\bigl(\frac{1}{4},\frac{1}{4},0\bigr)$ and fractional $\mathrm{ZZ}(\theta)\sim\mathrm{Can}\bigl(a,0,0\bigr)$ gates offers more promising ISA selections as they exhibit shorter gate duration, higher gate accuracy, and stronger synthesis capability.
The synthesis capability or computational power of basis gates can be geometrically illustrated by monodrome polytopes within the Weyl chamber. The coverage set for CX depicted in Figure 13 implies that
  1. (1)
    One $\mathrm{CX}$ gate is required to synthesize 2Q gates $\sim\mathrm{Can}\bigl(\frac{1}{2},0,0\bigr)$, i.e., $\mathrm{CX}$-equivalent gates $(V_{1}\otimes V_{2})\mathrm{CX}(V_{3}\otimes V_{4})$;
  2. (2)
    Two $\mathrm{CX}$ gates are required to synthesize 2Q gates $\sim\mathrm{Can}(a,b,0)$, i.e., $(V_{1}\otimes V_{2})\mathrm{CX}(V_{3}\otimes V_{4})\mathrm{CX}(V_{5}\otimes V_{6})$;
  3. (3)
    Three $\mathrm{CX}$ gates are required to synthesize 2Q gates $\sim\mathrm{Can}(a,b,c)$, i.e., $(V_{1}\otimes V_{2})\mathrm{CX}(V_{3}\otimes V_{4})\mathrm{CX}(V_{5}\otimes V_{6})\mathrm{CX}(V_{7}\otimes V_{8})$.
We assume the cost of one $\mathrm{CX}$ gate is $1.0$, polytopes in different colors denotes the minimal circuit cost (duration) for the coverage set if synthesized by $\mathrm{CX}$ and arbitrary 1Q gates. That is, on average, the number of $\mathrm{CX}$ gates required to synthesize arbitrary 2Q gates is $3$. In contrast, the number for SQiSW ISA is $2.21$ (huang2023quantum).
Monodromy polytope theory (peterson2020fixed) provides a framework for determining the synthesis coverage set and circuit cost (in 2Q depth) for any set of basis gates with specified costs, while the specific gate decomposition process is left to the synthesizer to complete. For the selected ISAs in Table 2 with the basis gate costs assumed in Equation 6, Figures 13, 13, 14, 15, 16 and 17 describes their coverage sets, respectively. With the enrichment of quantum ISA (e.g., combining gate families, involving mirror gates) and heterogeneous basis gate cost settings, the coverage set reveals a richer variety of convex polyhedra. That implies more optimization effects for the ISA-ware routing mechanism in Canopus.
A.3. 2Q gate mirroring
The mirror symmetry of a 2Q gate $U$ is defined as the composition of the original gate and a $\mathrm{SWAP}$ gate (proctor2022measuring), i.e., $\mathrm{SWAP}\cdot U$. For example, $\mathrm{CX}$ and $\mathrm{iSWAP}$ is a typical pair of mirror gates as shown below.
=              $^{\dagger} SS^{\dagger}$ $\mathrm{iSWAP}$ $\mathrm{iSWAP}$ $H$ $H$ $H$ $H$ $^{\dagger} SS^{\dagger}$
Figure 18. Morrir symmetry for $\mathrm{Can}(a,b,0)$ and $\mathrm{Can}(\frac{1}{2},b^{\prime},c^{\prime})$ gate families.
In general, the mirroring rule for Canonical coefficients is described as
[…]
The mirror pair of $\mathrm{CX}$ and $\mathrm{iSWAP}$ is a special case implying that a CX-iSWAP combination ISA could result in lower overhead in routing-synthesis collaborative optimization. yale2024noise once considers inserting $\mathrm{SWAP}$ gates to get mirrored gates with lower synthesis overhead compared to the original gates, given the all-to-all topology and continuous $\mathrm{ZZ}(\theta)$ gate set on ion-trapped hardware. mckinney2024mirage discusses that integrating $\mathrm{\sqrt{iSWAP}}$’s mirror gate, i.e., $\mathrm{ECP}\sim\mathrm{Can}\bigl(\frac{1}{4},\frac{1}{4},0\bigr)$ gate, into the powerful SQiSW ISA, could further improve the ISA’s synthesis capability and end-to-end routing-synthesis co-optimization on limited topologies.
Appendix B Commutative relation of canonical gates
Herein we present detailed proof for Theorem 1. The if direction is trivial, and hence we justify the only if direction, relying on the following two lemmas.
Lemma 0.
Let $A$, $B$ be two Hermitian matrices with eigenvalues in the range [2,2)[-2,2). If $e^{-i\frac{\pi}{2}A},e^{-i\frac{\pi}{2}B}]=0$ then $[A,B]=0$.
Proof.
This follows from the fact that compatible observables (commuting operators) can be simultaneously diagonalized. In this case, the respective unitary matrix $e^{-i\frac{\pi}{2}A}$ commutes with $e^{-i\frac{\pi}{2}B}$. Denote by $A_{\lambda}$ the eigenspace corresponding to the eigenvalue $\lambda$ of $e^{-i\frac{\pi}{2}A}$, i.e. $e^{-i\frac{\pi}{2}A}=\oplus_{\lambda}\lambda A_{\lambda}$. Then we have
(11) $\forall\vec{v}\in A_{\lambda},\,e^{-i\frac{\pi}{2}B}e^{-i\frac{\pi}{2}A}\vec{v}=e^{-i\frac{\pi}{2}B}\lambda\vec{v}=\lambda e^{-i\frac{\pi}{2}B}\vec{v}=e^{-i\frac{\pi}{2}A}e^{-i\frac{\pi}{2}B}\vec{v},$
and thus $e^{-i\frac{\pi}{2}B}\vec{v}\in A_{\lambda}$. Thus $A_{\lambda}$ is $e^{-i\frac{\pi}{2}B}$-invariant and the restriction $e^{-i\frac{\pi}{2}B}\big\rvert_{A_{\lambda}}$ of $e^{-i\frac{\pi}{2}B}$ to $A_{\lambda}$ is still unitary since it preserves inner products. Hence it is diagonalizable and we can find an orthonormal basis $w_{\lambda_{1}},w_{\lambda_{2}},\ldots,w_{\lambda_{k}}$ consisting of eigenvectors of $e^{-i\frac{\pi}{2}B}\big\rvert_{A_{\lambda}}$. Note that these are also eigenvectors of $e^{-i\frac{\pi}{2}A}$ (with eigenvalue $\lambda$). Following the same token as above, for each eigenspace $E_{\lambda_{i}}$ of $e^{-i\frac{\pi}{2}A}$, we can construct an orthonormal basis $\beta_{i}$ for it consisting of eigenvectors of $e^{-i\frac{\pi}{2}B}$. Finally since the eigenspaces of different eigenvalues of $e^{-i\frac{\pi}{2}A}$ are orthogonal to each other, $\beta=\cup_{i}\beta_{i}$ forms an orthonormal basis of the entire Hilbert space $\mathcal{H}_{n}$ consisting of the coeigenvectors of both $e^{-i\frac{\pi}{2}A}$ and $e^{-i\frac{\pi}{2}B}$.
Now let $U$ be a unitary matrix with the vectors in $\beta$ being its columns, then
(12) $\begin{aligned} U^{\dagger}e^{-i\frac{\pi}{2}A}U&=D_{A}\\ U^{\dagger}e^{-i\frac{\pi}{2}B}U&=D_{B}\end{aligned}$
In general, an eigenvector of $e^{-i\frac{\pi}{2}A}$ need not be that of $A$. However, since $A$ has its eigenvalues in the range [2,2)[-2,2), the map
(13) $f:[-2,2)\rightarrow U(1),a\rightarrow e^{-i\frac{\pi}{2}a}$
is injective. Consequently different eigenvalues of $A$ correspond to different eigenvalues of $e^{-i\frac{\pi}{2}A}$, and hence the eigenspaces of $e^{-i\frac{\pi}{2}A}$ and $A$ coincide. Therefore, we have that
(14) $\begin{aligned} U^{\dagger}AU&=\Sigma_{A}\\ U^{\dagger}BU&=\Sigma_{B}\end{aligned}$
and since $\Sigma_{A},\Sigma_{B}]=0$ as they are diagonal, $[A,B]=0$. We obtain the desired result.
Lemma 0.
Let $P_{1}=(a_{1}X_{1}X_{2}+b_{1}Y_{1}Y_{2}+c_{1}Z_{1}Z_{2})I_{3}$, $P_{2}=I_{1}(a_{2}X_{2}X_{3}+b_{2}Y_{2}Y_{3}+c_{2}Z_{2}Z_{3})$ with $|c_{1}|\leq b_{1}\leq a_{1}\leq\frac{1}{2}$, $|c_{2}|\leq b_{2}\leq a_{2}\leq\frac{1}{2}$. If $P_{1},P_{2}]=0$ and $P_{1},P_{2}\neq 0$, then $b_{1}=b_{2}=c_{1}=c_{2}=0$.
Proof.
Consider the product $P_{1}P_{2}$. We assume for the sake of contradiction that $b_{1}\neq 0$. Using $[X,Y]=2iZ$, $[Y,Z]=2iX$, $[Z,X]=2iY$, we expand
$[P_{1},P_{2}]$ $=2i\bigl(a_{1}b_{2}\,X_{1}Z_{2}Y_{3}-b_{1}a_{2}\,Y_{1}Z_{2}X_{3}+b_{1}c_{2}\,Y_{1}X_{2}Z_{3}\bigr)-2i\bigl(a_{1}c_{2}\,X_{1}Y_{2}Z_{3}+c_{1}a_{2}\,Z_{1}Y_{2}X_{3}+c_{1}b_{2}\,Z_{1}X_{2}Y_{3}\bigr).$
Since the each Pauli string is linearly independent in the $8\times 8$ operator basis, e.g. term $Y_{1}Z_{2}X_{3}$ cannot be canceled out by any other terms, contradictory to the fact that $P_{1},P_{2}]=0$. Hence, vanishing of $P_{1},P_{2}]$ requires
$a_{1}b_{2}=a_{1}c_{2}=b_{1}c_{2}=b_{1}a_{2}=c_{1}a_{2}=c_{1}b_{2}=0.$
Since $P_{1},P_{2}\neq 0$, at least $a_{1},a_{2}$ is nonzero, leading to $b_{1}=b_{2}=c_{1}=c_{2}=0$. ∎
Using Lemma 1 and Lemma 2 above, it is straightforward to prove Theorem 1. We see that $\|P_{1}\|\leq\|a_{1}X_{1}X_{2}I_{3}\|+\|b_{1}Y_{1}Y_{2}I_{3}\|+\|c_{1}Z_{1}Z_{2}I_{3}\|\leq|a_{1}|+|b_{1}|+|c_{1}|\leq\frac{3}{2}$, where $\|\cdot\|$ is the operator norm. Hence, eigenvalues of $P_{1}$ are in range of [2,2)[-2,2). Same as the eigenvalues of $P_{2}$. Now if $e^{-i\frac{\pi}{2}P_{1}},e^{-i\frac{\pi}{2}P_{2}}]=0$, then we have that $P_{1},P_{2}]=0$ according to Lemma 1, and thus $b_{1}=b_{2}=c_{1}=c_{2}=0$ according to Lemma 2, which proves the only if direction.
Appendix C Detailed evaluation results
Table 5. Routing overhead in terms of $C_{\mathrm{count}}$ for different compilers across different topologies and quantum ISAs.
Chain CX ISA ZZPhase ISA SQiSW ISA ZZPhase_ ISA SQiSW_ ISA Het ISA
Bench sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop
bigadd 2.53 2.44 1.59 1.92 2.35 2.26 1.53 1.97 2.39 2.23 1.42 1.98 1.90 1.81 1.57 1.59 1.94 1.85 1.49 1.57 1.95 1.85 1.31 1.77
bv 2.67 4.06 10.94 2.00 2.67 4.06 10.94 2.33 2.38 3.12 8.12 1.88 2.09 3.01 8.12 1.63 2.03 2.89 7.22 1.61 2.12 3.07 7.76 1.79
ising 1.00 1.00 1.00 1.00 0.38 0.38 0.38 0.38 0.75 0.75 0.75 0.75 0.38 0.38 0.38 0.38 0.75 0.75 0.75 0.75 0.38 0.38 0.38 0.38
knn 2.60 4.02 1.48 1.29 2.40 3.93 1.22 1.23 2.43 3.46 1.21 1.39 1.93 3.01 1.13 1.04 1.98 2.90 1.06 1.07 1.98 3.11 1.01 1.08
multi 2.32 4.97 2.53 2.68 2.18 4.83 2.28 2.49 2.26 4.17 1.99 2.38 1.79 3.67 1.84 2.02 1.81 3.56 1.73 2.04 1.83 3.78 1.68 2.01
qec9 4.44 12.34 6.88 3.56 4.44 12.34 5.33 3.53 3.89 9.52 3.47 2.84 3.43 9.05 5.23 2.77 3.25 8.45 3.98 2.56 3.52 9.34 4.27 2.77
qft 1.74 1.50 1.78 1.49 1.51 1.45 2.02 1.45 1.31 1.12 1.53 1.12 1.12 1.05 1.50 1.05 1.19 1.00 1.32 1.00 1.16 1.10 1.41 1.10
qpe 2.77 3.32 3.15 2.86 2.46 3.09 2.89 2.75 2.08 2.50 2.23 2.13 1.82 2.27 2.07 1.99 1.89 2.24 2.13 1.88 1.89 2.36 2.35 2.04
qram 2.94 5.37 2.75 3.23 2.75 5.21 2.73 3.02 2.63 4.44 2.53 2.80 2.16 3.93 2.29 2.45 2.19 3.80 2.22 2.37 2.22 4.06 2.26 2.43
sat 2.44 2.66 1.88 2.29 2.23 2.43 1.38 2.13 2.24 2.36 1.42 2.03 1.79 1.92 1.39 1.67 1.85 1.99 1.32 1.76 1.83 1.96 1.13 1.73
swapt 2.87 4.02 1.43 1.29 2.67 3.93 1.22 1.23 2.66 3.46 1.21 1.39 2.13 3.01 1.02 1.10 2.17 2.90 1.07 1.07 2.19 3.11 1.00 1.08
wstate 1.00 1.00 1.00 1.00 1.00 1.00 0.99 1.00 1.50 1.50 1.47 1.50 1.00 1.00 0.99 1.00 1.00 1.00 0.99 1.00 1.00 1.00 0.99 1.00
[…]
HHex CX ISA ZZPhase ISA SQiSW ISA ZZPhase_ ISA SQiSW_ ISA Het ISA
Bench sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop
bigadd 2.17 2.10 1.90 1.89 2.00 1.93 2.35 1.92 2.14 2.02 1.86 1.84 1.65 1.58 1.24 1.45 1.70 1.63 1.57 1.45 1.69 1.61 1.33 1.53
bv 3.28 2.22 7.22 1.94 3.28 2.22 7.23 1.83 3.00 2.12 5.83 1.50 2.58 1.80 5.51 1.51 2.47 1.75 4.94 1.42 2.64 1.82 5.35 1.68
ising 1.72 3.20 1.64 1.42 1.10 2.77 0.83 0.83 1.29 2.50 1.17 1.11 0.90 2.12 0.84 0.67 1.23 2.24 1.15 1.04 0.92 2.18 0.86 0.58
knn 2.18 2.57 2.17 1.49 1.98 2.43 1.81 1.39 2.12 2.33 1.91 1.54 1.63 1.92 1.61 1.16 1.70 1.93 1.66 1.25 1.66 1.97 1.75 1.17
multi 2.23 3.48 2.11 2.24 2.09 3.35 1.69 2.10 2.19 3.07 1.95 2.00 1.72 2.61 1.64 1.64 1.75 2.57 1.70 1.75 1.75 2.68 1.53 1.67
qec9 3.16 4.78 4.84 3.16 3.16 4.78 4.82 3.19 2.91 4.03 4.20 2.84 2.49 3.64 3.81 2.43 2.39 3.45 4.94 2.39 2.55 3.73 5.23 2.53
qft 1.91 2.62 2.44 1.67 1.60 2.35 1.83 1.52 1.43 1.97 1.61 1.27 1.19 1.73 1.50 1.12 1.31 1.78 1.46 1.12 1.24 1.80 1.59 1.16
qpe 2.58 2.90 2.89 2.43 2.15 2.59 2.44 2.10 1.94 2.18 2.30 1.86 1.61 1.92 1.96 1.62 1.77 1.97 1.86 1.68 1.67 1.99 1.82 1.67
qram 2.52 4.32 3.03 2.42 2.32 4.15 3.23 2.31 2.35 3.68 2.52 2.19 1.87 3.17 2.58 1.86 1.92 3.11 2.18 1.85 1.91 3.27 1.94 1.87
sat 2.27 2.29 1.60 2.00 2.05 2.06 1.28 1.81 2.10 2.10 1.35 1.83 1.65 1.66 1.33 1.44 1.74 1.74 1.26 1.52 1.69 1.69 1.03 1.49
swapt 2.24 2.57 2.06 1.50 2.04 2.43 1.78 1.42 2.18 2.33 1.98 1.56 1.68 1.92 1.63 1.14 1.74 1.93 2.20 1.21 1.71 1.97 1.82 1.18
wstate 2.65 2.04 2.60 1.69 2.65 2.04 2.47 1.46 2.71 2.28 2.21 1.67 2.19 1.75 2.07 1.50 2.10 1.69 1.78 1.35 2.23 1.78 1.97 1.60
[…]
Square CX ISA ZZPhase ISA SQiSW ISA ZZPhase_ ISA SQiSW_ ISA Het ISA
Bench sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop
bigadd 1.62 1.89 1.41 1.38 1.44 1.71 1.02 1.18 1.75 1.92 1.42 1.43 1.26 1.44 1.17 1.04 1.34 1.51 1.23 1.14 1.27 1.46 1.10 1.02
bv 2.72 2.22 4.39 1.50 2.72 2.22 4.39 1.50 2.75 2.38 3.67 1.50 2.23 1.87 3.44 1.22 2.14 1.81 2.94 1.31 2.28 1.90 3.17 1.22
ising 1.00 1.00 1.78 1.00 0.38 0.38 1.16 0.38 0.75 0.75 1.33 0.75 0.38 0.38 0.90 0.38 0.75 0.75 1.27 0.75 0.38 0.38 0.96 0.38
knn 1.79 2.64 1.88 1.29 1.57 2.43 1.52 1.23 1.88 2.46 1.59 1.42 1.35 1.96 1.40 1.04 1.45 2.01 1.46 1.07 1.38 2.00 1.28 1.05
multi 1.84 2.81 1.99 1.56 1.68 2.67 1.49 1.42 1.95 2.64 1.64 1.58 1.44 2.14 1.32 1.24 1.50 2.14 1.64 1.28 1.47 2.19 1.21 1.26
qec9 2.06 4.44 3.69 1.78 2.06 4.44 3.80 1.72 2.20 3.89 2.88 1.71 1.74 3.43 2.30 1.44 1.69 3.25 2.34 1.50 1.77 3.52 2.42 1.46
qft 1.41 2.30 2.09 1.35 1.03 1.88 1.41 1.05 1.06 1.75 1.50 1.00 0.79 1.42 1.08 0.80 0.99 1.60 1.37 0.94 0.82 1.47 1.06 0.81
qpe 1.68 2.50 2.06 1.46 1.22 2.00 1.42 1.18 1.26 1.91 1.47 1.11 0.95 1.53 1.08 0.93 1.18 1.74 1.48 0.99 0.98 1.58 1.26 0.90
qram 1.88 2.67 2.40 1.60 1.65 2.48 1.78 1.44 1.88 2.48 2.00 1.55 1.39 1.99 1.70 1.22 1.50 2.02 1.74 1.27 1.41 2.04 1.47 1.29
sat 1.65 2.09 1.44 1.54 1.42 1.87 1.22 1.34 1.70 2.06 1.25 1.45 1.22 1.55 1.11 1.11 1.34 1.64 1.14 1.21 1.24 1.58 0.94 1.12
swapt 1.75 2.64 2.06 1.29 1.54 2.43 1.14 1.18 1.85 2.46 1.46 1.44 1.33 1.96 1.38 1.04 1.43 2.01 1.47 1.11 1.35 2.00 1.33 1.07
wstate 1.00 1.00 1.25 1.00 1.00 1.00 1.39 1.00 1.50 1.50 1.82 1.50 1.00 1.00 1.53 1.00 1.00 1.00 1.31 1.00 1.00 1.00 1.32 1.00
[…]
Table 6. Routing overhead in terms of $C_{\mathrm{depth}}$ for different compilers across different topologies and quantum ISAs.
Chain CX ISA ZZPhase ISA SQiSW ISA ZZPhase_ ISA SQiSW_ ISA Het ISA
Bench sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop
bigadd 2.82 1.95 1.45 1.73 2.66 1.81 1.27 1.81 2.65 1.88 1.40 1.88 2.14 1.48 1.16 1.48 2.15 1.53 1.30 1.48 2.19 1.51 1.05 1.46
bv 2.83 2.72 5.06 2.11 2.83 2.72 5.06 2.39 2.62 2.12 4.12 2.12 2.26 2.01 3.73 1.74 2.19 1.89 3.58 1.72 2.29 2.07 3.99 1.85
ising 1.00 1.00 1.00 1.00 0.46 0.46 0.46 0.46 0.75 0.75 0.75 0.75 0.46 0.46 0.46 0.46 0.75 0.75 0.75 0.75 0.46 0.46 0.46 0.46
knn 3.16 2.66 1.65 1.39 2.90 2.53 1.45 1.31 2.76 2.32 1.35 1.35 2.27 1.95 1.32 1.05 2.32 1.95 1.17 1.10 2.33 2.03 1.12 1.11
multi 2.33 3.49 2.45 2.17 2.23 3.41 2.14 1.93 2.28 3.06 1.88 1.96 1.82 2.64 1.89 1.64 1.82 2.56 1.57 1.59 1.86 2.71 1.58 1.58
qec9 5.33 7.33 6.00 4.08 5.33 7.33 5.58 3.83 4.38 5.75 4.00 3.38 4.00 5.40 4.70 3.36 3.75 5.04 4.33 2.79 4.12 5.58 4.75 3.29
qft 2.85 1.50 2.50 1.50 1.95 1.43 3.02 1.42 2.14 1.12 2.24 1.12 1.53 1.04 2.27 1.03 2.01 1.01 2.02 1.01 1.58 1.08 2.29 1.08
qpe 4.00 2.97 3.66 2.83 3.32 2.69 3.67 2.77 3.00 2.23 2.73 2.12 2.49 1.99 2.62 1.83 2.75 2.01 2.64 1.90 2.58 2.06 2.70 1.71
qram 2.94 3.79 2.45 2.50 2.80 3.73 2.24 2.24 2.64 3.21 2.16 2.28 2.21 2.84 2.03 1.90 2.19 2.72 2.03 1.91 2.27 2.93 1.76 1.94
sat 2.28 2.00 1.77 1.88 2.12 1.85 1.25 1.64 2.19 1.87 1.30 1.75 1.73 1.50 1.36 1.37 1.77 1.54 1.20 1.48 1.77 1.52 1.06 1.38
swapt 3.42 2.66 1.58 1.39 3.15 2.53 1.44 1.31 2.98 2.32 1.37 1.35 2.45 1.95 1.12 1.13 2.50 1.95 1.16 1.10 2.52 2.03 1.11 1.11
wstate 1.00 1.00 1.04 1.00 1.00 1.00 1.02 1.00 1.50 1.50 1.50 1.50 1.00 1.00 1.02 1.00 1.00 1.00 1.02 1.00 1.00 1.00 1.01 1.00
[…]
HHex CX ISA ZZPhase ISA SQiSW ISA ZZPhase_ ISA SQiSW_ ISA Het ISA
Bench sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop
bigadd 2.55 1.77 1.81 1.80 2.39 1.62 2.12 1.80 2.46 1.78 1.59 1.69 1.95 1.37 1.32 1.32 1.97 1.43 1.56 1.35 1.99 1.38 1.19 1.34
bv 3.06 2.11 3.28 2.00 3.06 2.11 3.28 1.89 2.79 1.96 2.79 1.58 2.41 1.69 2.51 1.56 2.31 1.64 2.33 1.53 2.46 1.71 2.52 1.85
ising 4.50 6.25 4.25 3.00 3.58 5.92 2.71 2.12 3.38 5.25 2.44 2.62 2.75 4.54 2.42 1.60 3.12 4.50 2.88 2.62 2.83 4.67 2.44 1.19
knn 2.52 2.19 1.94 1.48 2.24 2.01 1.68 1.35 2.30 2.02 1.90 1.56 1.80 1.60 1.49 1.09 1.90 1.66 1.69 1.30 1.84 1.65 1.56 1.19
multi 2.14 2.76 1.88 1.92 2.05 2.69 1.50 1.80 2.11 2.51 1.82 1.93 1.68 2.12 1.57 1.54 1.68 2.06 1.59 1.75 1.71 2.17 1.41 1.57
qec9 4.50 3.83 4.25 3.25 4.50 3.83 4.29 3.42 4.06 3.38 5.25 3.06 3.51 2.96 3.29 2.64 3.33 2.83 5.38 2.75 3.60 3.02 5.77 3.08
qft 3.56 3.62 3.80 2.42 2.87 3.18 2.78 2.34 2.67 2.73 2.44 1.78 2.16 2.36 2.34 1.84 2.46 2.48 2.43 1.51 2.23 2.44 2.55 1.85
qpe 4.53 2.98 4.13 2.97 3.67 2.64 3.37 2.52 3.41 2.26 3.43 2.48 2.77 1.97 2.76 1.83 3.14 2.03 2.49 2.38 2.87 2.04 2.29 1.90
qram 2.86 3.01 2.87 2.17 2.67 2.84 2.44 1.92 2.66 2.71 2.16 2.04 2.15 2.25 2.26 1.74 2.17 2.24 2.02 1.62 2.20 2.30 2.02 1.80
sat 2.18 1.84 1.48 1.77 2.03 1.69 1.17 1.62 2.09 1.81 1.28 1.68 1.65 1.40 1.27 1.33 1.70 1.46 1.18 1.39 1.69 1.42 0.95 1.36
swapt 2.47 2.19 2.00 1.53 2.19 2.01 1.64 1.34 2.25 2.02 1.79 1.58 1.76 1.60 1.51 1.13 1.86 1.66 1.77 1.17 1.80 1.65 1.68 1.21
wstate 3.07 1.93 2.43 1.57 3.07 1.93 2.02 1.57 3.03 2.25 2.41 2.04 2.49 1.69 1.86 1.65 2.38 1.64 1.89 1.36 2.54 1.71 2.07 1.42
[…]
Square CX ISA ZZPhase ISA SQiSW ISA ZZPhase_ ISA SQiSW_ ISA Het ISA
Bench sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop sabre toqm bqskit canop
bigadd 1.90 1.48 1.28 1.33 1.74 1.32 1.04 1.16 2.00 1.60 1.18 1.46 1.49 1.16 1.01 1.04 1.55 1.23 1.14 1.09 1.51 1.17 0.98 1.01
bv 2.44 1.89 3.00 1.56 2.44 1.89 3.00 1.61 2.46 2.04 2.83 1.50 2.00 1.58 2.42 1.22 1.92 1.53 2.22 1.25 2.04 1.61 2.36 1.22
ising 1.00 1.00 8.00 1.00 0.46 0.46 6.75 0.46 0.75 0.75 4.31 0.75 0.46 0.46 2.67 0.46 0.75 0.75 3.62 0.75 0.46 0.46 3.56 0.46
knn 1.98 2.11 1.90 1.39 1.69 1.85 1.48 1.16 1.92 1.96 1.62 1.32 1.41 1.50 1.36 1.01 1.55 1.62 1.52 1.10 1.44 1.53 1.27 1.02
multi 1.86 1.98 1.94 1.57 1.73 1.92 1.48 1.43 1.98 1.99 1.52 1.71 1.48 1.58 1.30 1.28 1.52 1.56 1.55 1.23 1.51 1.61 1.18 1.23
qec9 3.00 3.58 3.67 1.92 3.00 3.58 3.67 1.83 3.06 3.38 3.31 1.81 2.47 2.85 2.71 1.51 2.38 2.71 2.75 1.50 2.52 2.92 2.87 1.62
qft 3.09 2.88 4.41 2.15 2.22 2.32 2.98 1.54 2.32 2.20 3.14 1.57 1.72 1.75 2.11 1.45 2.17 2.02 3.05 1.91 1.78 1.81 2.03 1.37
qpe 2.80 2.50 3.02 2.22 2.10 1.95 2.04 1.61 2.10 1.90 2.16 1.49 1.62 1.49 1.58 1.18 1.95 1.76 2.08 1.46 1.67 1.54 1.62 1.24
qram 2.06 1.83 2.35 1.35 1.88 1.71 1.58 1.28 2.03 1.86 1.83 1.55 1.56 1.42 1.56 1.28 1.63 1.46 1.71 1.21 1.59 1.45 1.34 1.23
sat 1.57 1.55 1.44 1.50 1.41 1.42 1.16 1.32 1.68 1.68 1.22 1.45 1.22 1.22 1.09 1.06 1.30 1.28 1.11 1.20 1.24 1.24 0.86 1.12
swapt 1.97 2.11 1.98 1.39 1.68 1.85 1.23 1.24 1.94 1.96 1.49 1.31 1.41 1.50 1.27 1.05 1.55 1.62 1.54 1.08 1.44 1.53 1.22 1.00
wstate 1.00 1.00 1.36 1.00 1.00 1.00 1.23 1.00 1.50 1.50 1.79 1.50 1.00 1.00 1.85 1.00 1.00 1.00 1.29 1.00 1.00 1.00 1.32 1.00
[…]